OpenStudy (ronaldo7):

7x+2= x√5 Can anyone walk me through this please

OpenStudy (anonymous):

First, you need to gather all the terms containing x on one side of the equation. Can you do that?

OpenStudy (ronaldo7):

yeah

OpenStudy (anonymous):

Go ahead. What do you have?

OpenStudy (ronaldo7):

so 7x-x√5 = -2

OpenStudy (anonymous):

Terrific. Now, x is a common factor on the left hand side. Can you factor out the x?

OpenStudy (ronaldo7):

yeah x(7-√5) =-2

OpenStudy (anonymous):

Great. Now, to solve for x, divide both sides by \(7-\sqrt{5}\)

OpenStudy (ronaldo7):

ok so it becomes x = -2/7-√5 and thats it?

OpenStudy (anonymous):

Well not quite. Depends on what your teacher expects. Most times we don't want to leave a square root (irrational number) in the denominator, so we would rationalize the denominator. Do you know what I'm talking about?

OpenStudy (ronaldo7):

oh yeah you have to multiply by conjugate

OpenStudy (ronaldo7):

i just remembered

OpenStudy (anonymous):

Exactly. Multiply your answer by the conjugate over the conjugate.

OpenStudy (ronaldo7):

so it becomes -1/27

OpenStudy (ronaldo7):

right?

OpenStudy (anonymous):

Don't think so. Let me detail it a bit...

OpenStudy (ronaldo7):

ok

OpenStudy (anonymous):

\[x=\frac{ -2 }{ 7-\sqrt{5} }\times \frac{ 7+\sqrt{5} }{7+\sqrt{5} }\]Can you complete this?

OpenStudy (ronaldo7):

Oh i forget to multiply the top i did it in my head

OpenStudy (ronaldo7):

\[-14+2√5 / 49 +5\]

OpenStudy (anonymous):

Close. Check the signs in the numerator. And the denominator is going to be 49-5

OpenStudy (ronaldo7):

yeah is se

OpenStudy (ronaldo7):

see*

OpenStudy (ronaldo7):

So the final answer is\[−7+1√5/22\]

OpenStudy (ronaldo7):

i simplified

OpenStudy (anonymous):

Excellent. And you don't need to write the 1 before the square root sign.

OpenStudy (ronaldo7):

But hold on i have the same exact answer on my test and I got it wrong i put a plus or minus but my teacher took off a ll the points he usually takes of like half

OpenStudy (ronaldo7):

and h crossed out my entire process

OpenStudy (ronaldo7):

i squared both sides when i started solving is that why?

OpenStudy (ronaldo7):

can i not do that

OpenStudy (anonymous):

Can't help you there. You'll have to ask your teacher about that. But the work you've done here is correct.

OpenStudy (ronaldo7):

Ok no problem thank you so much for the help

OpenStudy (anonymous):

You wouldn't solve this problem by squaring. IN this case, you can't get rid of the square root by squaring both sides.

OpenStudy (anonymous):

You're welcome